OpenStudy (anonymous):

Find the value of y so that the line passing through (6, y) and (7, 5) has a slope of -3.

OpenStudy (johnweldon1993):

Well we know slope is \[\large \frac{y_2 - y_1}{x_2 - x_1}\] You have your points... (6 , y) (x1 , y1) and (7 , 5) (x2 , y2) So \[\large \frac{5 - y}{7 - 6}\] \[\large \frac{5 - y}{1}\] What does y have to be for this to = -3 ?

OpenStudy (anonymous):

what that guy said...

OpenStudy (anonymous):

Not suree....

OpenStudy (anonymous):

read again

OpenStudy (johnweldon1993):

@Autobee7128 If you see that last fraction I came down to... \[\large \frac{5 - y}{1} = -3\] We know that anything divided by 1 is that number...so we dont even need that \[\large 5 - y = -3\] What does y equal?